OpenStudy (anonymous):

describe two different ways to solve -10=1/4(8y-12)

OpenStudy (anonymous):

Cross Multiply... :)

OpenStudy (anonymous):

Is this your question? \[-10 = \frac{1}{4}(8y-12) \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

First Method: Multiply by 4 both the sides first and tell me what did you get..

OpenStudy (anonymous):

multiply 4 times 8?

OpenStudy (anonymous):

No: Like this: \[-10 \times 4 = \frac{1}{4}(8y-12) \times 4 \implies -40 = 8y - 12 \implies -40 + 12 = 8y\]

OpenStudy (anonymous):

oh ok i get it now thanks alot :)

OpenStudy (anonymous):

Can you find y from there?? What did you get?

OpenStudy (anonymous):

-3.5